OpenStudy (anonymous):

Find the Domain and range of each function. 1) \[y=\sqrt{x-1}+2\] 2) \[y=-\sqrt{x+3}-1\] 3) \[y=1/2\sqrt{x}+3\] 4) \[y=-\sqrt{x+4}-1\] Also, how would I graph these?

OpenStudy (anonymous):

The domain of the first function is [1, infinity), and its range is [2, infinity).

OpenStudy (anonymous):

infinity as in greater than or equal to?

OpenStudy (anonymous):

What do you mean? [1, infinity) means all the numbers from 1 to infinity, inclusive.

OpenStudy (mertsj):

Do you understand that the radicand cannot be negative?

OpenStudy (anonymous):

Listen to Mertsj: He knows what he is talking about. :P

OpenStudy (mertsj):

And answer him.

OpenStudy (anonymous):

The inifinity thing doesnt seem familiar, i haven't heard about that in class. the radicand is the number under the radical, yes

OpenStudy (mertsj):

\[x-1\ge0\] \[x \ge1\]

OpenStudy (mertsj):

So since the radicand, which is x-1, must be greater than or equal to 1, x can be any real number that is greater than or equal to 1. And that is the domain. Lots of times instead of saying greater than or equal to 1, mathematicians use interval notation which would be:

OpenStudy (mertsj):

\[[1,\infty)\]

OpenStudy (mertsj):

The bracket shows that x could be 1 but not infinity, sisnce there is no largest number.

OpenStudy (mertsj):

Or you could use set builder notation and say that the domain is:

OpenStudy (mertsj):

\[{x|x \ge 1}\]

OpenStudy (mertsj):

Do we need to discuss the range of the first one?

OpenStudy (mertsj):

The smallest value of sqrt(x-1) is 0 so the smallest y value is 2. So the range is all real numbers greater than or equal to 2

OpenStudy (anonymous):

so is x−1≥0 okay for the domain? I think my teacher's okay with the simply one.

OpenStudy (anonymous):

I'll try.... what do you need help with?

OpenStudy (anonymous):

@dpaInc finding the domain and range

OpenStudy (anonymous):

@dumbcow can you help

OpenStudy (anonymous):

@Luis Rivera @.Sam. @saifoo.khan @KingGeorge

OpenStudy (kinggeorge):

Which one are we working on right now?

OpenStudy (anonymous):

the first one since i wasnt even sure of the answer

OpenStudy (kinggeorge):

First, lets start with the domain. Since the domain of a function \(g(x)=\sqrt{x}\) is \(x \geq 0\), we know that the domain of a function \(f(x)=\sqrt{x-1}\) is \(x-1 \geq 0\). In other words, \(x\geq1\) is the domain. Since your original function is \[f(x) = \sqrt{x-1} +2\]Your domain is \(x\geq1\). Does this make sense so far?

OpenStudy (anonymous):

so you pay attention to what's under the radical?

OpenStudy (kinggeorge):

For the domain, yes. Since a function \(f(x) =x+2\) is defined everywhere, we can ignore the \(+2\) in your function if we're finding the domain.

OpenStudy (anonymous):

what about the range?

OpenStudy (kinggeorge):

For the range, we know that the function is always increasing, so we only have to find the smallest value possible, and extend it to infinity. Since \(\sqrt{x}\geq 0\) we know that \(\sqrt{x-1}\geq0\) so if we add 2, the range of \(y=\sqrt{x-1}+2\), is \(y\geq2\)

OpenStudy (anonymous):

how would the graph of that function look like? like a sketch

OpenStudy (kinggeorge):